# Applications of higher order Markov models and Pressure Index to strategize controlled run chases in Twenty20 cricket

### v1

@@ 6.1 Estimating the order of Markov chain
    Higher order dependencies: A higher-order Markov model captures more complex dependencies by considering a longer history of past states. This means an order-3 model takes into account the previous three states to predict the current state, whereas a lower-order model considers fewer past states.
@@
    Uniformity in model structure: Using the same order for all sequences ensures that the model structure is consistent across all sequences. This uniformity is crucial for combining data from different sequences while constructing a single transition matrix.
@@
    Avoiding underfitting: Using a lower order might underfit the data, missing out on important dependencies captured by higher order models.
Using the highest order ensures that the full extent of dependencies present in the sequences is captured, providing a more accurate and informative transition matrix. While this might be more complex, it avoids the risk of missing critical patterns that could have been obscured by using a lower order model.
6.2 Applying the Markov model on Pressure Index
From each sequence of the entire data, every possible state, say $x_{t}$, is considered in the columns as the current state and every set of preceding states in the form of a tuple $(x_{t-3},x_{t-2},x_{t-1})$ is considered in the rows as the previous states. Note that $\{X_{n}:n\geq 1\}$ is the sequence of Pressure Index values at the end of each over during the second innings of a match, and so, $x_{t}\in S$, where $S$ is the state space that contains all possible values the Pressure Index of a team can take. The transition probabilities $p_{ij}$ are then obtained from the data and the resulting matrix satisfies the properties mentioned in Definition 4.1.2. The obtained transition matrix  [^5] The transition matrix obtained using PI values of 3378 Twenty20 matches is of order 49803 $\times$ 1053. exhibits the set of all possible probabilities with which a $PI$ value (current state) succeeds a $3$-tuple of $PI$ values (previous state).
It is important to note that the system inevitably produces some negative $PI$ values, as it is designed to take into account the calculations until the end of the innings. In matches where the team batting second achieves its target, the final $PI$ value comes out to be negative. We will simply avoid such values for practical purposes since calculating the Pressure Index after a team has won the match does not make sense.
After the transition matrix has been obtained, any new sequence consisting of $s$ entries ($s\geq 3$) can be inserted to predict the outcome of the unknown $(s+1)$th entry of the sequence. The prediction process is not very simple. While a tuple of previous states is inserted, the model works in a way such that it searches for occurrences of the exact set of previous states (maintaining the order) in the transition matrix. If multiple current states are observed corresponding to a set of previous states with nonzero probabilities in the transition matrix, then the expected value of the current state is considered as the final outcome. However, if no occurrences of current states are found with nonzero probability for a set of previous states, then the sum of the states in the tuple is calculated and sets of previous states whose entries add up to the same sum are searched in the transition matrix. The expected value of the current state is then calculated following similar steps as described earlier. In either case, an outcome is obtained irrespective of whether the exact previous states are present in the observed dataset.
The following example shows the outcome when $(1.49,1.63,1.48)$ is entered as the new sequence. In cricketing terms, this translates to a situation where the Pressure Index of a team batting second have been recorded as $1.49$, $1.63$, and $1.48$ at the end of three consecutive overs during the second innings of a Twenty20 match. Given the target and the number of overs the innings is into, we are interested in predicting the probable Pressure Index of the team at the end of the next over.
The device provides a list of all possible Pressure Index values a team has exhibited in the training sample of $3378$ Twenty20 matches after having $PI$ values $1.49$, $1.63$, $1.48$ in that order. The probabilities corresponding to the occurrences of each possible $PI$ value are listed from the transition matrix, and then finally the expected outcome of the Pressure Index value is calculated from these probabilities.
Pressure Index Probability of occurrence Pressure Index Probability of occurrence
0.00 0.010310341 1.66 0.041241365
0.31 0.010310341 1.67 0.030931024
0.66 0.010310341 1.68 0.041241365
0.82 0.010310341 1.69 0.020620683
0.88 0.010310341 1.70 0.015465512
1.21 0.010310341 1.71 0.030931024
1.25 0.010310341 1.72 0.046396536
1.36 0.010310341 1.73 0.046396536
1.39 0.010310341 1.74 0.025775853
1.40 0.010310341 1.75 0.030931024
1.42 0.005155171 1.76 0.056706877
1.43 0.010310341 1.79 0.005155171
1.46 0.010310341 1.80 0.010310341
1.47 0.010310341 1.81 0.030931024
1.48 0.010310341 1.82 0.020620683
1.51 0.015465512 1.84 0.030931024
1.52 0.015465512 1.86 0.041241365
1.54 0.015465512 1.87 0.010310341
1.56 0.005155171 1.90 0.010310341
1.57 0.010310341 1.91 0.010310341
1.58 0.015465512 1.94 0.003402413
1.59 0.036086194 1.95 0.005155171
1.60 0.018867925 2.00 0.010310341
1.61 0.020620683 2.02 0.020620683
1.62 0.020620683 2.23 0.010310341
1.64 0.034333436 2.39 0.010310341
Expected Pressure Index outcome: 1.576342
Table 6.2.1 Expected outcomes of $PI$ values based on previous states
This means that if a team faces a situation where they have recorded Pressure Index values of $1.49$, $1.63$, $1.48$ in that order at the end of three consecutive overs during the second innings of a Twenty20 match, then the expected Pressure Index value at the end of the next over would be approximately $1.576342$.
6.3 Strategy for a controlled run chase using the Markov model
It has been observed from extensive studies that if the Pressure Curve during a run chase reaches values less than $0.5$, then the match is most likely to be won by the team batting second. From [Saikia et al., 2019], it is understood that
and for the different subintervals of $PI$ values between $[0.5,3.5]$, viz. $[0.5,\lambda_{i_{1}}]$, $[\lambda_{i_{1}},\lambda_{i_{2}}]$, $\dots$, $[\lambda_{i_{m}},3.5]$, the values of
generally start with $1$ but keep on decreasing until they eventually reach 00. An immediate strategy for teams batting second in Twenty20 matches would be to try to bring down the Pressure Index as low as possible, and effectively below $0.5$, at some stage during the innings in order to guarantee a win.
With the transition matrix already obtained and the target known at the start of the innings, the expected $PI$ value at the end of the next over at any stage during the run chase can be calculated. This allows teams batting second to strategize accordingly and set mini targets for the next few overs to try to reduce their Pressure Index to below $0.5$.
For example, in the situation elaborated in Table 6.2.1, for the occurrence of a sequence of $PI$ values $1.49$, $1.63$, $1.48$, the expected Pressure Index at the end of the next over is $1.576342$ and the probability of this Pressure Index being less than $0.5$ is only about $2.0620682\%$. This is a situation where it is visibly difficult for the team batting second to immediately get a hold on the match.
(expected $\mathbb{P}(\text{expected }PI<0.5)=\mathbb{P}(PI=0)+\mathbb{P}(PI=0.31)=0.020620682$ (15)
6.3.1 Case Study: Using the model in a T20I match
Consider the Twenty20 match played between Pakistan and West Indies at National Stadium, Karachi on April 03, 2018. Pakistan were chasing a target of 154 and won the match by 8 wickets. The Pressure Index values of Pakistan’s batting innings are as follows.
Overs Cum runs Cum wickets Pressure Index Expected PI
1 5 0 1.038638 -
2 15 0 1.043422 -
3 32 0 0.990569 -
4 51 0 0.908084 1.090972
5 57 0 0.933522 0.823888
6 62 1 1.035470 0.816428
7 71 1 1.029399 0.919259
8 79 1 1.032575 0.941079
9 86 1 1.048083 1.052562
10 89 1 1.131503 1.13
11 98 1 1.113852 1.082961
12 107 1 1.084413 1.09
13 114 2 1.147530 1.1407
14 120 2 1.173178 1.177427
15 132 2 0.940734 0.947136
16 139 2 0.829457 0.826328
17 154 2 0 0.0734211
Table 6.3.1.1 Pressure Index of Pakistan at the end of every over
Note that the expected PI values have gradually converged towards the actual PI values, especially after the end of the 9th over. The strategy of determining the PI values after every over thus works well in this run chase.
6.3.2 A counter example for the strategy
Consider the Twenty20 match played between Bangladesh and Pakistan at Sher-E Bangla National Stadium, Mirpur on April 24, 2015. The hosts were chasing a target of 142 and won the match by 7 wickets. The Pressure Index values of Bangladesh’s batting innings are listed in the table below.
Overs Cum runs Cum wickets Pressure Index Expected PI
1 14 1 1.041062 -
2 16 2 1.081729 -
3 25 2 1.194013 -
4 31 2 1.224512 1.150613
5 34 3 1.294495 1.226884
6 44 3 1.371937 1.2118
7 50 3 1.414531 1.363112
8 54 3 1.495481 1.464337
9 65 3 1.458339 1.6028
10 75 3 1.429256 1.430688
11 89 3 1.286981 1.285
12 97 3 1.261276 1.265888
13 107 3 1.152069 1.153505
14 116 3 1.027661 1.03
15 123 3 0.929088 0.929653
16 138 3 0.252506 0.964372
17 143 3 0 0.2066667
Table 6.3.2.1 Pressure Index of Bangladesh at the end of every over
Bangladesh’s expected PI was gradually converging towards the actual PI values until it predicted $0.964372$ at the end of the 16th over, deviating significantly from the actual PI $0.252506$ which happened on the brink of a 15-run over. The probability of PI being less than $0.5$ at that stage was only $0.111152\%$. Incidentally, this was also the highest number of runs scored in an over among all 47 overs in that match. This brought down their $PI$ to much lesser than $0.5$ and they completed the run chase in the next over itself. This example shows how the indices can be monitored throughout the innings and how continuously revising the probabilities of keeping the Pressure Index within a suitable range helps a team remain in control during a run chase in Twenty20 cricket.
6.4 Building a strategy using phase-wise Markov models
Although 6.3 provides an insight into how Pressure Index values can be monitored throughout an innings to help build a strategy for a controlled run chase, it emerges as quite a general approach irrespective of a team’s specific patterns during run chases in certain situations. For a better understanding, one might be interested in studying the patterns of where a team’s Pressure Index lies at different stages during a successful run chase in Twenty20 cricket.
The $PI$ values at the end of the 6th, 10th and 16th overs of teams during several successful run chases (comprising 2773 out of the 3378 matches from our initial training set) are considered. The choice of these specific stages corresponds to the importance of the end of the powerplay (6th over), the halfway mark of the innings (10th over), and the start of the death overs (17th over). AIC, BIC and Kolmogorov-Smirnov tests have been performed over a few known distributions to determine which known distribution fits the data corresponding to $PI$ values the best.
Criteria Gamma Exponential Weibull
$AIC$ $-1513.574$ $7094.485$ $-1065.882$
$BIC$ $-1501.721$ $7100.412$ $-1054.029$
K-S test statistic $(D)$ $0.022463$ $0.50916$ $0.074801$
K-S p-value $0.1221$ $2.2\times 10^{-16}$ $6.905\times 10^{-14}$
Table 6.4.1 Results for $PI$ values at the end of the 6th over
Criteria Gamma Exponential Weibull
$AIC$ $1285.369$ $8229.105$ $1511.110$
$BIC$ $1297.222$ $8235.032$ $1522.962$
K-S test statistic $(D)$ $0.017787$ $0.47144$ $0.062958$
K-S p-value $0.345$ $2.2\times 10^{-16}$ $5.858\times 10^{-10}$
Table 6.4.2 Results for $PI$ values at the end of the 10th over
Criteria Gamma Exponential Weibull
$AIC$ $4931.218$ $9351.203$ $5006.789$
$BIC$ $4943.068$ $9357.128$ $5018.64$
K-S test statistic $(D)$ $0.025706$ $0.38125$ $0.036553$
K-S p-value $0.05169$ $2.2\times 10^{-16}$ $0.001233$
Table 6.4.3 Results for $PI$ values at the end of the 16th over
Lower AIC, BIC, and K-S test statistic values, along with higher p-values, denote that the gamma distribution fits well for the $PI$ values corresponding to each section of the data. This gives us an idea of how the $PI$ values at three different stages of an innings are distributed.
Figure 2: Density plots of phase-wise Pressure Index
Phases of the innings Estimated shape parameter Estimated rate parameter
Powerplay $68.72744$ $57.66806$
Middle overs $18.66357$ $11.17790$
Death overs $3.926073$ $1.985266$
Table 6.4.4 — Estimated parameters of Gamma distribution
Separate transition matrices corresponding to the three different stages of an innings can be constructed based on the process described in 6.2. Outcomes resulting from a tuple consisting of three $PI$ values will then be obtained based on the transition matrix corresponding to the specific section of the innings only, irrespective of $PI$ values occurring in the other two sections. Once the estimated $PI$ value of a team for a certain over of the innings is obtained, one can attempt to bring it down to $0.5$ as early as possible, according to the theory described in 6.3. The three-phase partition of the innings avoids the overgeneralization of teams’ strategies in run chases.
7 Direction of future research
Tracing Pressure curves ([Bhattacharjee and Lemmer, 2016]) and revising Pressure Index values with the help of the proposed Markov model helps to build a strategy to keep teams batting second in Twenty20 matches in control throughout the chase. However, several potential developments to the model can be considered as part of future research:
  1. i.
    Different teams have different patterns of a run chase. For example, teams with greater batting depths might score at a faster rate even after the fall of a wicket and still succeed, but the same strategy might not work for teams with lesser batting depths. Therefore, errors are expected to occur when using the transition probabilities corresponding to one team to predict Pressure Index values for a team with a different composition.
  2. ii.
    The proposed Markov model does not take into account any factors regarding venues, which might play a significant role in run chases in Twenty20 matches. Modifying the model to incorporate the effects of different venues might help to increase the accuracy when predicting Pressure Index values.
  3. iii.
    The model could be improved by considering the differences in the qualities and strengths of the teams involved in a match. Apart from venues, the ability of teams winning close contests, and the strength of opposition bowlers and teams, might affect a team’s performance. For example, a Pressure Index of say, $1.42$, at some stage of the innings against a stronger opposition may have a different impact compared to that against a weaker opposition. In matches between a top-ranked team and a lower-ranked team, the former is expected to capitalize towards the end of the match even from less favorable situations, primarily due to the significant difference in experience possessed by the two teams.
@@ 7 Direction of future research
    The formula (5.1.1) for Pressure Index restricts its usage to the second innings only, primarily because of the behavior of the Pressure curve in case of unsuccessful run chases. Certain variance stabilizing transformations (suggested by Saikia, Bhattacharjee and Mukherjee, $2019$), where $\frac{CRRR}{IRRR}$ is raised to a power $(\frac{CRRR}{IRRR})^{\alpha}$ for some suitable choice of $\alpha$, help to restrict the abrupt rise in $PI$ values at later stages of the innings during unsuccessful run chases. Such transformations can be applied in constructing a similar Markov model if one intends to use it for prediction purposes for teams batting first as well.

### v2

@@ 8 Direction of future research
    Different teams have different patterns of a run chase. For example, teams with greater batting depths might score at a faster rate even after the fall of a wicket and still succeed but the same strategy might not work for teams with lesser batting depths. Therefore, errors are expected to occur when using the transition probabilities corresponding to one team to predict Pressure Index values for a team with a different composition.
@@
    Besides the effects of contextual factors like home ground advantage and tournament-specific models discussed in Section 5.4, external factors like venues, psychological factors like player-player coalition or player-coach coalition etc. might also play key roles in run chases in Twenty20 matches. Modification of the model to incorporate such external factors, using small sample methodologies, might increase the accuracy while predicting Pressure Index values.
@@
    The model could be improved by considering the differences in the qualities and strengths of the teams involved in a match. Apart from the external factors, the ability of teams winning close contests, strengths of oppositions etc. might affect a team’s performance. For example, a Pressure Index of say, $1.42$, at some stage of the innings against a stronger opposition may have a different impact compared to that against a weaker opposition. In matches between a top-ranked team and a lower-ranked team, the former is expected to capitalize towards the end of the match even from less favorable situations, primarily due to the significant difference in experience possessed by the two teams.
@@
    The formula in 6 restricts the usage of Pressure Index to second innings only. Certain variance stabilizing transformations (suggested by [17]), where $\frac{CRRR}{IRRR}$ is raised to a power $(\frac{CRRR}{IRRR})^{\alpha}$ for some suitable choice of $\alpha$, help to restrict the abrupt rise in $PI$ values at later stages of the innings during unsuccessful run chases. Such transformations can be applied in constructing a similar Markov model if one intends to use it for prediction purposes for teams batting first as well.
@@
Appendix A Appendix
A.1 Necessary tables
Table 24: Predictive performance of the Markov model for some tournaments/series
Competition MAE RMSE Coverage Markov Usage
Tournaments T20 Internationals 1.7749 4.0912 82.43% 93.64%
Abu Dhabi T20 League 7.1149 13.9755 82.35% 94.12%
Africa Regional T20 Cup 6.2063 16.5576 82.55% 95.22%
Bangabandhu T20 Cup 3.1105 4.7792 87.96% 95.24%
Bangladesh Premier League 1.0662 3.5498 80.58% 96.12%
BBL 1.8664 3.1107 84.43% 95.92%
CPL 5.5460 8.9030 84.31% 95.10%
CSA T20 Challenge 3.3286 4.4902 84.16% 95.52%
Dhaka T20 League 4.8326 6.5975 81.73% 94.31%
Emerging Teams Asia Cup 1.9141 3.7764 84.05% 93.25%
ILT20 1.0259 2.7928 83.86% 95.48%
IPL 0.8885 7.8697 84.16% 95.86%
Lanka Premier League 3.9791 9.9558 84.62% 95.78%
Major League Cricket 2.0930 5.1130 83.61% 95.48%
National T20 Cup 5.5214 8.1287 84.90% 95.41%
PSL 0.6081 3.6784 84.53% 95.93%
Ram Slam T20 Challenge 2.1653 5.0489 81.12% 96.76%
SA20 League 6.3580 13.7701 84.16% 93.70%
SLC Twenty-20 Tournament 4.3099 9.4595 82.05% 96.46%
Super Smash 1.5733 2.8123 84.49% 95.06%
Super T20 Provincial Tournament 5.429 12.769 92.80% 93.07%
Syed Mushtaq Ali Trophy 1.6174 3.3080 84.97% 95.66%
T20 Blast 6.4472 14.1590 83.37% 95.90%
Few randomly picked series Afghanistan A in Oman T20s 7.8297 5.8597 78.79% 96.97%
Australia v Sri Lanka 2022 0.4093 1.1497 84.29% 92.86%
BBL 2023 0.7681 2.7041 87.59% 97.19%
CPL 2023 4.1487 3.0939 84.75% 94.92%
CPL 2024 3.9750 4.8647 84.25% 95.00%
ICC T20 WC Warm-up 2022 1.3045 3.0810 85.02% 87.50%
Ireland A in Bangladesh T20Is 0.2406 0.4900 100.00% 100.00%
New Zealand v England 2019 1.0183 6.7465 88.24% 88.24%
SA20 League 2024 5.3580 7.7701 78.16% 93.70%
SL2020 4.0565 6.1665 79.49% 94.12%
South Africa A in Zimbabwe T20Is 1.0273 4.6591 83.33% 89.58%
Syed Mushtaq Ali Trophy 2018–19 0.9208 1.9607 83.78% 99.10%
West Indies A in Nepal T20s 1.1591 5.0552 84.67% 87.65%
A.2 Pseudo codes for useful Algorithms
Algorithm 2 Determine optimal Markov chain order
1:Dataset of $M$ matches with PI sequences
2:Maximum order $k_{max}$
3:Optimal order $k^{*}$
4:Split data into 80% training and 20% test sets
5:for $k=1$ to $k_{max}$ do
6:  Initialize transition count matrix $N^{(k)}$
7:  $N_{transitions}\leftarrow\sum_{m=1}^{M}(T_{m}-k)$ $\triangleright$ Equation 12
8:  for each match $m$ in training set do
9:   for $t=k+1$ to $T_{m}$ do
10:     Extract previous state: $(s_{t-k},\ldots,s_{t-1})$
11:     Extract current state: $s_{t}$
12:     Increment $N^{(k)}[(s_{t-k},\ldots,s_{t-1}),s_{t}]$
13:   end for
14:  end for
15:  Calculate $m_{k}$ = number of unique transitions
16:  Calculate log-likelihood: $L_{k}\leftarrow\sum_{i=1}^{N_{transitions}}\log P(s_{t}^{(i)}|\text{previous state}^{(i)})$
17:  Calculate $AIC_{k}\leftarrow 2m_{k}-2L_{k}$ $\triangleright$ Equation 15
18:  Calculate $BIC_{k}\leftarrow m_{k}\log(N_{transitions})-2L_{k}$ $\triangleright$ Equation 16
19:  Calculate parameter-to-observation ratio: $\rho_{k}\leftarrow\frac{m_{k}}{N_{transitions}}$
20:end for
21:Select $k^{*}\leftarrow\arg\min_{k}\{AIC_{k},BIC_{k}:\rho_{k}<1\}$ return $k^{*}$
Algorithm 3 Discretize Pressure Index values
1:Continuous PI values $\{PI_{1},PI_{2},\ldots,PI_{n}\}$
2:Precision level $\delta$
3:Discretized PI values $\{PI^{\prime}_{1},PI^{\prime}_{2},\ldots,PI^{\prime}_{n}\}$
4:for each $PI_{i}$ in dataset do
5:  if $PI_{i}<0$ then
6:   $PI^{\prime}_{i}\leftarrow 0$ $\triangleright$ Censoring negative values
7:  else
8:   $PI^{\prime}_{i}\leftarrow\delta\times\lfloor\frac{PI_{i}}{\delta}+0.5\rfloor$ $\triangleright$ Round to nearest $\delta$
9:  end if
10:end forreturn $\{PI^{\prime}_{1},PI^{\prime}_{2},\ldots,PI^{\prime}_{n}\}$
Algorithm 4 Predict next Pressure Index value
1:Transition matrix $\mathcal{T}$
2:Previous states $(s_{t-k},\ldots,s_{t-1})$
3:Phase-specific Gamma parameters $(\alpha,\beta)$
4:Confidence level $\gamma$
5:Predicted PI value $\hat{s}_{t}$ with confidence interval $\hat{s}_{t}^{lower},\hat{s}_{t}^{upper}]$
6:if $(s_{t-k},\ldots,s_{t-1})$ exists in $\mathcal{T}$ then
7:  Extract all possible next states: $\{s_{t}^{(1)},s_{t}^{(2)},\ldots,s_{t}^{(n)}\}$
8:  Extract corresponding probabilities: $\{p_{1},p_{2},\ldots,p_{n}\}$
9:  Calculate expected value: $\hat{s}_{t}\leftarrow\sum_{i=1}^{n}p_{i}\cdot s_{t}^{(i)}$
10:else
11:  Calculate sum: $S\leftarrow\sum_{j=t-k}^{t-1}s_{j}$
12:  Search for states in $\mathcal{T}$ where $\sum_{j=t-k}^{t-1}s_{j}^{\prime}=S$
13:  if matching states found then
14:   Calculate expected value from matching transitions
15:  else
16:   Use Gamma distribution fallback: $\hat{s}_{t}\leftarrow\frac{\alpha}{\beta}$ $\triangleright$ Mean of Gamma
17:  end if
18:end if
19:Calculate confidence interval using Gamma quantiles:
[…]
Algorithm 5 Fit phase-wise distributions to Pressure Index
1:PI values for phase $\phi\in\{\text{Powerplay, Middle, Death}\}$
2:Distribution families $\mathcal{D}=\{\text{Gamma, Exponential, Weibull}\}$
3:Best-fit distribution parameters for phase $\phi$
4:for each distribution $d\in\mathcal{D}$ do
5:  Estimate parameters $\theta_{d}$ using Maximum Likelihood Estimation
6:  Calculate log-likelihood: $L_{d}$
7:  Calculate $AIC_{d}\leftarrow 2|\theta_{d}|-2L_{d}$
8:  Calculate $BIC_{d}\leftarrow|\theta_{d}|\log(n)-2L_{d}$
9:  Perform Kolmogorov-Smirnov test: $(D_{d},p_{d})\leftarrow\text{KS-test}(PI,d,\theta_{d})$
10:end for
11:Select $d^{*}\leftarrow\arg\min_{d}\{AIC_{d}\}$ subject to $p_{d^{*}}>0.05$ return Distribution $d^{*}$ with parameters $\theta_{d^{*}}$
Algorithm 6 Parametric Bootstrap for Kolmogorov-Smirnov Goodness-of-Fit
1:Observed data $x_{1},\dots,x_{n}$; distribution family $\mathcal{D}$ (e.g., Gamma)
2:K–S statistic $D_{\text{obs}}$ and bootstrap p-value $\hat{p}$
3:Fit $\mathcal{D}$ to observed data to obtain parameter estimate $\hat{\theta}$
[…]
5:Choose bootstrap size $B$
6:for $b=1$ to $B$ do
[…]
8:  Fit $\mathcal{D}$ to bootstrap sample to obtain $\hat{\theta}^{(b)}$
9:  Compute $D^{(b)}=\sup_{x}|F_{n}^{(b)}(x)-F(x;\hat{\theta}^{(b)})|$
10:end for
11:Compute bootstrap p-value: $\hat{p}=\frac{1}{B}\sum_{b=1}^{B}\mathbf{1}\{D^{(b)}\geq D_{\text{obs}}\}$ return $D_{\text{obs}},\hat{p}$
Algorithm 7 Calculate Pressure Index at any stage during a run chase
1:Target $T$
2:Total balls $B$
3:Runs scored $R^{\prime}$
4:Balls faced $B^{\prime}$
5:Wickets fallen $w_{1},w_{2},\ldots,w_{k}$
6:Resources used RURU
7:Pressure Index PIPI
8:Calculate Initial Required Run Rate: $IRRR\leftarrow\frac{T\times 6}{B}$
9:Calculate Current Required Run Rate: $CRRR\leftarrow\frac{(T-R^{\prime})\times 6}{B-B^{\prime}}$
10:Initialize wicket weight sum: $\sum w_{i}\leftarrow 0$
11:for each fallen wicket at position $i$ do
12:  $\sum w_{i}\leftarrow\sum w_{i}+w_{i}$ $\triangleright$ Use weights from Table 1
13:end for
14:Calculate: $PI\leftarrow\frac{CRRR}{IRRR}\times\frac{1}{2}\left[e^{\frac{RU}{100}}+e^{\frac{\sum w_{i}}{11}}\right]$ return PIPI
Algorithm 8 Generate over-by-over strategic recommendations
1:Current over $t$
2:PI history $(PI_{1},\ldots,PI_{t})$
3:Target $T$
4:Wickets lost $w$
5:Home/Away status
6:Trained phase-wise models $\{\mathcal{M}_{\text{PP}},\mathcal{M}_{\text{MO}},\mathcal{M}_{\text{DO}}\}$
7:Strategic recommendation and target PI range
8:Determine current phase $\phi$ based on over $t$
9:Select appropriate model $\mathcal{M}_{\phi}$ and thresholds from Table 19
10:Predict next PI: $\hat{PI}_{t+1}\leftarrow\mathcal{M}_{\phi}(PI_{t-2},PI_{t-1},PI_{t})$
11:if $\hat{PI}_{t+1}\in[0,0.5)$ then
12:  Recommendation $\leftarrow$ ”Target Zone - Maintain aggressive scoring”
13:else if $\hat{PI}_{t+1}\in[0.5,1.5)$ then
14:  Recommendation $\leftarrow$ ”Acceptable Zone - Continue current approach”
15:else if $\hat{PI}_{t+1}\in[1.5,2.5)$ AND phase = Death Overs then
16:  Recommendation $\leftarrow$ ”Acceptable/Risky - Accelerate carefully”
17:else
18:  Recommendation $\leftarrow$ ”Avoid Zone - High risk, need immediate acceleration”
19:  Calculate required run rate adjustment
20:end if
21:Display: Current PI, Predicted PI range, Recommendation, Required runs per over return Recommendation with actionable targets